OpenStudy (anonymous):

Dr. Oguz counted the number of cubs in 10 tiger cub litters and then arranged the numbers from smallest to largest. The numbers of cubs in each litter were 1, 1, 2, 2, 3, 3, 3, 4, 5, and 6. The mean of these measurements is 3. What is the mean absolute deviation? A. 1.2 B. 1.5 C. 2 D. 12

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Do not post questions with no work shown. You should calculate the Absolute Deviation from 3 for each item.

OpenStudy (anonymous):

@tkhunny I don't know how

OpenStudy (anonymous):

Let's see...I think it's C...But I'm probably wrong..

OpenStudy (tkhunny):

Hogwash. |1-3| = 2 |2-3| = 1 |2-3| = 1 |3-3| = 0 |3-3| = 0 |3-3| = 0 |4-3| = 1 |5-3| = 2 |6-3| = 3 You can't do that?

OpenStudy (anonymous):

@tkhunny I didnt know it was that easy

OpenStudy (anonymous):

@tkhunny Thanks

OpenStudy (tkhunny):

I knew you could do it. Now add them up and divide by however many there are. Mean, Absolute Deviation. Done.

OpenStudy (anonymous):

10

OpenStudy (tkhunny):

That's the TOTAL of Absolute Deviations. Now the MEAN Absolute Deviation...

OpenStudy (anonymous):

How do you get it @tkhunny

OpenStudy (tkhunny):

Asked and answered. Mean --- Average --- Divide by something...

OpenStudy (anonymous):

3/10 is 0.3

OpenStudy (tkhunny):

We already know the MEAN of the data. That's 3. We already added things up and divided by 9. This gave 3, the Mean or Average. We need the Mean Absolute Deviation. Find the total of the absolute deviations. You did that. 10. Now divide by what to get the MEAN absolute deviation?